OpenStudy (jaylelile):

Use the information below to find EF such that ABC ~ DEF. AB = 36, BC = 24, DE = 48, B= 110 angle E = 110 angle, EF = ? A.32 B.36 C.68 D.72

OpenStudy (campbell_st):

well if the triangles are similar then use the ratio of corresponding sides to find EF so its \[\frac{DE}{AB} = \frac{EF}{BC}\] then \[\frac{48}{36} = \frac{EF}{24}\] hope it helps